I have know Idea what I am doing.I am trying to learn how to multiply and write in scientific notation. I have 3 problems I need to learn how to do.
(8.3*10to the negative 5th)(2*10^12)=
(7*10to the neg 3rd)(2.5*10^3)=
and
(2.1*10tothe neg8th)(2.5*10^3)=

Answer by scott8148(5892)

when you multiply in scientific notation, you multiply the coefficients (regular numbers) and add the powers of 10

(8.3*10^-5)(2*10^12)=(8.3*2)*10^(-5+12)=16.6*10^7
__ remember, only one number to the left of the decimal, so 1.66*10^8
__ the coefficient got smaller so the exponent had to get larger to keep the value the same

the other two problems work the same way